# for the Baryon Scattering (BaSc) Collaboration

### v1

@@ II.2.1 Estimation of priors for the conspiracy model
$m_{eff}(t)$ $=\ln\left(\frac{C(t)}{C(t+1)}\right)\,,$
$A_{eff}(t)$ $=e^{+m_{eff}(t)\,t}\ C(t)\,,$
@@
$\tilde{E}_{0}$ $=m_{eff}(t_{\rm ref})\times\mathcal{N}(1,0.1)\,,$
@@
For the parameters of the NN correlation functions, the priors of the overlap factors are estimated in the same way as described above. Since we model the NN energies as in Eq. (11), we need estimates for the $\delta_{nm}$ parameters. For the ground state interaction energy, we use the effective mass of the ratio correlation function at $t_{\rm ref}$ to estimate the prior mean and we choose to give it a 100% width
$\tilde{\delta}_{00}=m_{eff}^{R}(t_{\rm ref})\times\mathcal{N}(1,1)\,,$ (15)
@@
$\tilde{\delta}_{mn}=\tilde{\delta}_{nm}=\mathcal{N}(0,m_{eff}^{R}(t_{\rm ref}))\,.$ (16)
@@ II.3.1 Correlator analysis and the NN spectrum
Figure 2: Comparison of the ground state NN energy and energy shift in the $T_{1g}$ channel using the conspiracy (C) and agnostic (A) models (top). For the former, we show the preferred final model while for the latter, we vary the number of exponentials used to model the single ($n$) and NN ($nn$) correlators. The band and dashed lines are from the conspiracy model with 3-states for the single nucleon, to guide the eye. The relative probability for a given model ($w_{\rm model})$ at each $t_{\rm min}$ is determined from the Bayes Factor. We also show stability of the conspiracy model with respect to the number of states used to describe the single nucleon correlators (bottom). The $Q$-value factor is similar to the frequentist $P$-value but uses the augmented $\chi^{2}/{\rm dof}$ Lepage et al. (2002).
@@
    Fits with three or more states for the NN correlator are all very consistent, while those with only 2 states are consistent at the 1-2 sigma level but systematically different for all $t_{\rm min}$;
@@
In addition, we will show that with the conspiracy model, the resulting values and uncertainties of these $\delta_{00}$ values are more stable with respect to variation of the number of exponentials, the value of $t_{\rm min}^{N,NN}$ and other user choices.
@@
Overall, for all these variations, we find that the conspiracy model provides a very stable extraction of the ground-state NN energy and interaction energy. We have shown these stability plots for the lowest level in the $T_{1g}$ irrep, and we find similar stability for all irreps considered in this work. In Fig. 3, we plot the stability of ${\delta}_{00}$ in all 15 irreps/levels considered in this work for two of these systematics, and the remainder are plotted in Appendix A.
Figure 3: Stability of the ground state interaction energy, ${\delta}_{00}$, for all 15 irreps/levels used in this work for the deuteron. Note that the two levels in $A_{2}(1)$ have overlapping values of ${\delta}_{00}$.
@@ II.3.2 Amplitude analysis
Given a spectrum, we turn to applying the single-channel QC2 for which there is a one-to-one mapping between each energy level and a value of the amplitude. With our increased precision, the results have become sensitive to two sources of systematic uncertainty as compared to Ref. Hörz et al. (2021):
@@
    The final NN energy is sensitive to whether it is taken from the fit, or reconstructed using ${\delta}_{00}$ and the continuum dispersion relation of the single nucleons;
@@
    In some of the boosted channels, our results have become sensitive to $S-D$ mixing.
@@
It is expected that the leading discretization effects to the spectrum cancel in the energy splitting, ${\delta}_{00}=E_{NN}-E_{1}-E_{2}$, suggesting that using the continuum dispersion relation to reconstruct the single nucleon energies $E_{i}=\sqrt{M_{N}^{2}+n_{i}(2\pi/L)^{2}}$ is more appropriate for the QC2 as its derivation assumed such a dispersion relation. This expectation was recently formally derived in Ref. Hansen and Peterken (2024).
Regarding the $S-D$ partial wave mixing, Ref. Briceño et al. (2013) discussed in detail the expected mixing in these channels. Depending upon the total momentum boost vector, different linear combinations of the spectrum results from different irreps can be used to remove the leading dependence upon the $S-D$ mixing, such as
@@
$\frac{q\cot{\delta}}{m}=\frac{-1}{am}+\frac{1}{2}(mr)\left(\frac{q}{m}\right)^{2}+(m{\lambda})^{3}\left(\frac{q}{m}\right)^{4}+\cdots$
we determine the scattering length ($a$) and effective range ($r$). The scattering length is very stable under variations of the ERE order, while the effective range shows sensitivity to the truncation. Some of the analysis results are presented in Table 3.
Table 3: Results of the ERE parameters in the deuteron channel at $\mathcal{O}(q_{\rm cm}^{2})$ and $\mathcal{O}(q_{\rm cm}^{4})$ in the ERE expansion.
@@
Figure 4: Resulting phase shift expressed as $q\cot{\delta}$ (top) and ${\delta}$ in degrees (bottom) from the single-channel QC2 analysis after irrep averaging. The solid points are used in the analysis. The open data are the original results before irrep averaging. The dotted line in the $q\cot{\delta}$ plot is $\sqrt{-q^{2}}$, i.e. the virtual bound state condition. A crossing between the $q\cot\delta$ and this curve implies the presence of a virtual bound state. The NN ONLINE - PWA phase shift is obtained from http://nn-online.org using the Partial Wave Analysis of experimental data Stoks et al. (1993).
@@ II.4 Di-neutron channel results
We perform a similar stability analysis for the di-neutron scattering channel, finding the same optimal model used for the deuteron channel works well. In Fig. 5, we show the summary stability plots for the lab-frame energy shifts, ${\delta}_{00}$ for the 9 irreps/levels used in this work. Performing an amplitude analysis as with the deuteron, we determine the scattering length and effective range parameters at different orders in the ERE, listed in Table 4. This analysis also rules out a bound state in the di-neutron channel, and reveals the existence of a virtual bound state. For the di-neutron, the issue of partial wave mixing does not affect the analysis, as it does for the deuteron channel. Fig. 6 shows the resulting $q\cot{\delta}$ and phase shift. One interesting observation is that the effective range, $r_{0}$, in the di-neutron system is larger than the deuteron.
Figure 5: Stability of the ground state interaction energy , ${\delta}_{00}$ for all 8 irreps/levels used in this work in the di-neutron channel.
Table 4: Results of the ERE parameters in the dineutron channel at $\mathcal{O}(q_{\rm cm}^{2})$ and $\mathcal{O}(q_{\rm cm}^{4})$ in the ERE expansion.
@@ III.1 Adding HX operators to the basis
The results in the previous section did not include HX operators in the basis when constructing $C_{ij}(t)$. NPLQCD has suggested such operators may more efficiently couple to bound states than those where the nucleons are individually projected in momentum space Amarasinghe et al. (2023). Therefore, in this section, we explore the impact of including HX operators in the resulting identification of the NN spectrum.
The local HX operator we use is constructed in a similar manner and is designed to match that used by NPLQCD Beane et al. (2013a); Wagman et al. (2017); Amarasinghe et al. (2023) as closely as possible. For this study, we focus on $\bm{P}_{\rm tot}=0$ for both the deuteron in the the $T_{1g}$ irrep and the di-neutron in the $A_{1g}$ irrep. In the $T_{1g}$ symmetry channel, nine operators are used: $1$ HX and $8$ nucleon-nucleon, all of which are single-site operators. The correlator matrix is averaged over the three rows of the $T_{1g}$ irrep. In the $A_{1g}$ symmetry channel, five operators are used: $1$ HX and $4$ nucleon-nucleon. Again, the HX and individual nucleons are single-site operators. These operators are summarized in Table 5.
Table 5: Example $NN$ operators used in the $I=0\ T_{1g}$ and the $I=1\ A_{1g}$ channels for total momentum $\bm{P}=0$. The operator notation and the integers in the 2nd and 3rd columns are described in the text.
$NN$ Operator $I=0\ T_{1g}$ $I=1\ A_{1g}$
@@
(a) Hexaquark
(b) $N(0)N(0)$
(c) $N(1)N(1)$
Figure 10: Overlap factors for the HX and two momentum based operators in the isosinglet $T_{1g}$ channel.
(a) Level 1
(b) Level 1
(c) Level 1
Figure 11: Overlap factors for the HX and two momentum based operators in the isotriplet $A_{1g}$ channel.
@@ III.2 Off diagonal correlators with HX operators
NPLQCD has suggested that including local HX-type operators may be important for revealing their elusive bound state in the spectrum, arguing that the off-diagonal correlators may provide a more effective means to find them as discussed in Chapter 15 of Ref. Tews et al. (2022), as well as Ref. Amarasinghe et al. (2023). These authors use a toy model to illustrate their point. They introduce a three-state system with energies $E_{0}<E_{1}<E_{2}$ and two operators $A,B$, where the energies and overlap factors are given by
@@
Effective mass plots using values $\epsilon=0.05,\ \eta=0.3,\ \Delta=0.1,\ \delta=0.1$ are shown in the left plot of Fig. 12. The effective masses for the so-called principal correlators $\lambda_{0},\lambda_{1}$ are shown, as well as the effective mass corresponding to the off-diagonal correlator $C_{01}$. The principal correlators are the eigenvalues of the matrix $C(t_{0})^{-1/2}\,C(t)\,C(t_{0})^{-1/2}$. Here, we used $t_{0}=5$ with eigenvector pinning to identify the levels. The black curve shows the perfect extraction of $E_{0}$ starting at zero temporal separation from the off-diagonal correlator. The blue curve associated with principal correlator $\lambda_{0}(t)$ eventually plateaus at $E_{0}$, but not until very large temporal separations are reached due to the small coupling to the $E_{0}$ state. The red curve is associated with the principal correlator $\lambda_{1}(t)$ which plateaus eventually to $E_{1}$.
This is a very contrived and unrealistic (if not impossible to attain) toy model since it has been engineered to produce a perfect energy extraction from the off-diagonal correlator due to the exact zeros in the $Z_{A},Z_{B}$ factors. To achieve this situation in general would require an exact diagonalization of the full volume of lattice operators, not the use of a single operator. Making only a single change, allowing an arbitrarily small, non-zero overlap with these states,
$Z_{A}=(\epsilon_{1},\sqrt{1-\epsilon_{1}^{2}-\epsilon_{2}^{2}},\epsilon_{2}),\quad Z_{B}=(\epsilon_{2},\epsilon_{1},\sqrt{1-\epsilon_{1}^{2}-\epsilon_{2}^{2}}).$ (24)
immediately makes the effect go away, with the correlation function now scaling as,
@@
An even more realistic model is studied in the right-hand plot of Fig. 12. For this plot, two very simple changes are made from the original NPLQCD toy model: the overlaps with the ground state are not set to be the same for the states created by operators $A$ and $B$, and the exact zeros in the $Z_{A},Z_{B}$ factors are replaced by small but non-zero values:
@@
This model is still very unrealistic since there are no couplings to the many higher levels as there would be in actual lattice QCD computations. Lastly, consider a more realistic toy model with 10 levels and 3 operators $A,B,C$ and parameters chosen simply to be
[…]
@@
The effective masses corresponding to the three principal correlators $\lambda_{0}(t),\lambda_{1}(t),\lambda_{2}(t)$, as well as the effective mass corresponding to the off-diagonal correlator $C_{01}(t)$, are shown in Fig. 13. The principal correlators have been determined using $t_{0}=5$ in the left plot, and $t_{0}=25$ in the right plot. The off-diagonal correlator shows no advantage.
Figure 13: (Left) Effective masses corresponding to the three principal correlators $\lambda_{0}(t),\lambda_{1}(t),\lambda_{2}(t)$, shown as blue, red and green curves, respectively, and the effective mass corresponding to the off-diagonal correlator $C_{01}(t)$, shown as a black curve, for the 10-level, 3-operator toy model described in the text, with parameter values given in Eq. (27). Principal correlators have been computed using $t_{0}=5$. (Right) Same as the plot on the left, except the principal correlators have been determined using $t_{0}=25$. No advantage is observed for the off-diagonal correlator.
Our conclusion is that it is nearly impossible for an off-diagonal correlator to reveal a level hidden from the correlator matrix method, especially for a hexaquark operator that has been shown to contain copious amounts of excited contamination. In fact, negative weights in the spectral representation of an off-diagonal correlator with a very slow fall-off of contributions from higher-lying states are more likely to cause plateau misidentifications.
@@
Figure 14: (Top) Comparison of scattering phase shifts in the deuteron channel determined in Ref. Hörz et al. (2021), shown as red, blue, magenta points and magenta and gray bands, with phase shifts determined in Refs. Beane et al. (2013a, b), shown as green points. (Bottom left) Effective mass, shown as blue points, from Fig. 2 of Ref. Beane et al. (2017) for the deuteron on a $48^{3}$ spatial lattice from an off-diagonal correlator involving a local hexaquark operator at the source and a nucleon-nucleon operator of zero momentum at the sink. The horizontal green line indicates the approximate energy extraction from Fig. 4 of Ref. Beane et al. (2013a), with the violet line indicating approximately where the energy extraction from Ref. Hörz et al. (2021) would be for this $48^{3}$ lattice. (Bottom right) Effective mass, similar to that in the center plot, but for the dineutron.

### v2

@@ II.2.1 Estimation of priors for the conspiracy model
$m_{\rm eff}(t)$ $=\ln\left(\frac{C(t)}{C(t+1)}\right)\,$,
$A_{\rm eff}(t)$ $=e^{+m_{\rm eff}(t)\,t}\ C(t)\,$,
@@
$\tilde{E}_{0}$ $=m_{\rm eff}(t_{\rm ref})\times\mathcal{N}(1,0.1)\,$,
@@
For the parameters of the NN correlation functions, the priors of the overlap factors are estimated in the same way as described above. Since we model the NN energies as in Eq. (12), we need estimates for the $\delta E_{nm}$ parameters. For the ground state interaction energy, we use the effective mass of the ratio correlation function at $t_{\rm ref}$ to estimate the prior mean and we choose to give it a 100% width
$\tilde{\delta E}_{00}=m_{\rm eff}^{R}(t_{\rm ref})\times\mathcal{N}(1,1)\,$, (16)
@@
$\tilde{\delta E}_{mn}=\tilde{\delta E}_{nm}=\mathcal{N}(0,m_{\rm eff}^{R}(t_{\rm ref}))\,.$ (17)
@@
II.2.2 Comparison with other analysis strategies
A comparison with other analysis strategies in the literature provides some of the reasons for our use of this conspiracy model. The analysis procedures used in Refs. Fleming (2004); Fleming et al. (2009); Beane et al. (2009); Aubin and Orginos (2011a, b); Cushman and Fleming (2019); Romiti and Simula (2019); Cushman and Fleming (2020); Fischer et al. (2020); Fleming (2023) are based on the Prony Method de Prony (1795), the Generalized Pencil of Function (GPoF) method Hua and Sarkar (1989); Sarkar and Pereira (1995), or some combination of these two methods. A recent interest has arisen in using the Lanczos algorithm with such methods Wagman (2025); Hackett and Wagman (2025) which have been shown to be equivalent to the Prony-like methods Ostmeyer et al. (2025); Abbott et al. (2025). There has also been a suggestion of using Laplace filters Portelli and Tsang (2025). Common to all of these strategies is an attempt to isolate individual eigenvalues of the correlation functions and model their contributions by single-exponential forms. The large number of eigenstates which contribute to the spectral representation of any correlation function usually makes such isolation difficult and prone to unknown systematic errors. As with multi-exponential fits to any correlation functions, especially if the single nucleons are constructed with only simple 3-quark interpolating operators, the higher-lying exponential decay rates extracted should not be interpreted as stationary-state eigenvalues of QCD as the operator basis is not sufficient to reliably isolate these individual eigenstates. In all of these strategies, the higher-lying decay rates should be viewed as just nuisance parameters used to model the excited-state contamination, which is a sum of an infinite number of exponentials, by a sum of a finite (usually small) number of exponentials.
Another important point to keep in mind is the fact that the conspiracy model is applied after the use of the GEVP, which is also utilized to removed excited state contamination arising from the lower-lying NN interacting modes. The observed remaining excited state contamination comes from higher lying modes, $n\geq N_{\rm op}$ in Eq. (4). These other methods will also have to contend with this excited state contamination, which as we show below, most likely arises from excited states of the single nucleons.
Rather than trying to isolate an eigenstate that can be described by a single exponential, the conspiracy model is designed to capture the apparent physics contained in the NN correlators after the GEVP is used to diagonalize the basis of operators: the diagonal NN correlators are well described by Eq. (11) with small corrections induced by the interactions of these states. The conspiracy model not only allows for a systematic way to account for this excited-state contamination and its impact on the extraction of the ground-state energies, but also takes advantage of the positive-definite nature of these diagonal correlators to utilize regions in Euclidean time where the signal is not lost to the noise. In future work, it will be interesting to quantitatively compare the conspiracy model with the other analysis strategies mentioned above.
@@ II.3.1 Correlator analysis and the NN spectrum
Figure 2: Comparison of the ground state NN energy and energy shift in the $T_{1g}$ channel using the conspiracy (C) and agnostic (A) models (top). For the former, we show the preferred final model while for the latter, we vary the number of exponentials used to model the single ($n$) and NN (nnnn) correlators. The band and dashed lines are from the conspiracy model with 3-states for the single nucleon, to guide the eye. The relative probability for a given model ($w_{\rm model})$ at each $t_{\rm min}$ is determined from the Bayes Factor. We also show stability of the conspiracy model with respect to the number of states used to describe the single nucleon correlators (bottom). The $Q$-value factor is similar to the frequentist $P$-value but uses the augmented $\chi^{2}/{\rm dof}$ Lepage et al. (2002).
@@
    Fits with three or more states for the NN correlator are all consistent, while those with only 2 states are consistent at the 1-2 sigma level but systematically different for all $t_{\rm min}$;
@@
In addition, we will show that with the conspiracy model, the resulting values and uncertainties of these $\delta E_{00}$ values are more stable with respect to variation of the number of exponentials, the value of $t_{\rm min}^{N,NN}$ and other user choices.
@@
Overall, for all these variations, we find that the conspiracy model provides a stable extraction of the ground-state NN energy and interaction energy. We have shown these stability plots for the lowest level in the $T_{1g}$ irrep, and we find similar stability for all irreps considered in this work. In Fig. 3, we plot the stability of ${\delta}E_{00}$ in all 15 irreps/levels considered in this work for two of these systematics, and the remainder are plotted in Appendix A.
Figure 3: Stability of the ground state interaction energy, ${\delta}E_{00}$, for all 15 irreps/levels used in this work for the deuteron. Note that the two levels in $A_{2}(1)$ have overlapping values of ${\delta}E_{00}$.
@@ II.3.2 Amplitude analysis
Given a spectrum, we turn to applying the single-channel QC2 for which there is a one-to-one mapping between each energy level and a value of the amplitude. As in Ref. Hörz et al. (2021), we use the “spectrum method”, as described in Ref. Morningstar et al. (2017), to determine parameters of $q\cot{\delta}$ from the resulting energy levels. This is done by constructing the $\chi^{2}$ loss function in terms of $q^{2}_{\rm cm}$. With our increased precision, the results have become sensitive to two sources of systematic uncertainty as compared to Ref. Hörz et al. (2021):
@@
    The final NN energy is sensitive to whether it is taken from the fit, or reconstructed using ${\delta}E_{00}$ and the continuum dispersion relation of the single nucleons;
@@
    In some of the boosted channels, our results have become sensitive to SDS-D mixing.
@@
It is expected that the leading discretization effects to the spectrum cancel in the energy splitting, ${\delta}E_{00}=E_{NN}-E_{1}-E_{2}$, suggesting that using the continuum dispersion relation to reconstruct the single nucleon energies $E_{i}=\sqrt{M_{N}^{2}+n_{i}(2\pi/L)^{2}}$ is more appropriate for the QC2 as its derivation assumed such a dispersion relation. This expectation was recently formally derived in Ref. Hansen and Peterken (2024).
Regarding the SDS-D partial wave mixing, Ref. Briceño et al. (2013) discussed in detail the expected mixing in these channels using the QC2 formalism with spin 1/2 particles Briceno et al. (2013); Briceno (2014). Depending upon the total momentum boost vector, different linear combinations of the spectrum results from different irreps can be used to remove the leading dependence upon the SDS-D mixing, such as
@@
$\frac{q\cot{\delta}}{m_{\pi}}=\frac{-1}{am_{\pi}}+\frac{rm_{\pi}}{2}\left(\frac{q}{m_{\pi}}\right)^{2}+{\lambda}m_{\pi}^{3}\left(\frac{q}{m_{\pi}}\right)^{4}+\cdots$ (22)
we determine the scattering length ($a$) and effective range ($r$). The scattering length is stable under variations of the ERE order, while the effective range shows sensitivity to the truncation. Some of the analysis results are presented in Table 3.
Table 3: Results of the ERE parameters in the deuteron channel at $\mathcal{O}(q_{\rm c.m.}^{2})$ and $\mathcal{O}(q_{\rm c.m.}^{4})$ in the ERE expansion.
@@
Figure 4: Resulting phase shift expressed as $q\cot{\delta}$ (top) and $\delta$} in degrees (bottom) from the single-channel QC2 analysis after irrep averaging. The solid points are used in the analysis. The open data are the original results before irrep averaging. The dotted line in the $q\cot{\delta}$ plot is $\sqrt{-q^{2}}$, i.e. the virtual bound state condition. A crossing between the $q\cot\delta$ and this curve implies the presence of a virtual bound state. The NN ONLINE - PWA phase shift is obtained from http://$n$-online.org using the Partial Wave Analysis of experimental data Stoks et al. (1993).
@@ II.4 Di-neutron channel results
We perform a similar stability analysis for the di-neutron scattering channel, finding the same optimal model used for the deuteron channel works well. In Fig. 5, we show the summary stability plots for the lab-frame energy shifts, ${\delta}E_{00}$ for the eight irreps/levels used in this work. Performing an amplitude analysis as with the deuteron, we determine the scattering length and effective range parameters at different orders in the ERE, listed in Table 4. This analysis also rules out a bound state in the di-neutron channel, and reveals the existence of a virtual bound state. For the di-neutron, the issue of partial wave mixing does not affect the analysis, as it does for the deuteron channel. Fig. 6 shows the resulting $q\cot{\delta}$ and phase shift. One interesting observation is that the effective range, $r_{0}$, in the di-neutron system is larger than the deuteron.
Figure 5: Stability of the ground state interaction energy , ${\delta}E_{00}$ for all 8 irreps/levels used in this work in the di-neutron channel.
Table 4: Results of the ERE parameters in the dineutron channel at $\mathcal{O}(q_{\rm c.m.}^{2})$ and $\mathcal{O}(q_{\rm c.m.}^{4})$ in the ERE expansion.
@@ III.1 Adding HX operators to the basis
The results in the previous section did not include HX operators in the basis when constructing $C_{ij}(t)$. The authors of Ref. Amarasinghe et al. (2023) suggested such operators may more efficiently couple to bound states than those where the nucleons are individually projected in momentum space. Therefore, in this section, we explore the impact of including HX operators in the resulting identification of the NN spectrum.
The local HX operator we use is constructed in a similar manner and is designed to match that used in Refs. Beane et al. (2013a); Wagman et al. (2017); Amarasinghe et al. (2023) as closely as possible. For this study, we focus on $\bm{P}_{\rm tot}=0$ for both the deuteron in the the $T_{1g}$ irrep and the di-neutron in the $A_{1g}$ irrep. In the $T_{1g}$ symmetry channel, nine operators are used: $1$ HX and $8$ nucleon-nucleon, all of which are single-site operators. The correlator matrix is averaged over the three rows of the $T_{1g}$ irrep. In the $A_{1g}$ symmetry channel, five operators are used: $1$ HX and $4$ nucleon-nucleon. Again, the HX and individual nucleons are single-site operators. These operators are summarized in Table 5.
Table 5: Example NNNN operators used in the $I=0\ T_{1g}$ and the $I=1\ A_{1g}$ channels for total momentum $\bm{P}=0$. The operator notation and the integers in the 2nd and 3rd columns are described in the text.
NNNN Operator $I=0\ T_{1g}$ $I=1\ A_{1g}$
@@
Figure 10: Overlap factors for the hexaquark and two momentum based operators in the isosinglet $T_{1g}$ channel. $N(\bm{d}^{2})N(\bm{d}^{2})$ refers to an operator expected to predominantly create or annihilate a two-nucleon state in which one nucleon has momentum $\bm{p}=(2\pi/L)\bm{d}$, where $\bm{d}$ is a three-vector of integers, and the other nucleon has momentum $\bm{p}$.
Figure 11: Overlap factors for the hexaquark and two momentum-based NNNN operators in the isotriplet $A_{1g}$ channel (similar to Fig. 10).
@@ III.2 Off diagonal correlators with HX operators
In Chapter 15 of Ref. Tews et al. (2022) and in Ref. Amarasinghe et al. (2023), the authors suggested that including local HX-type operators may be important for revealing their elusive bound state in the spectrum, arguing that the off-diagonal correlators may provide a more effective means to find them. These authors use a toy model to illustrate their point. They introduce a three-state system with energies $E_{0}<E_{1}<E_{2}$ and two operators $A,B$, where the energies and overlap factors are given by
@@
Effective mass plots using values $\epsilon=0.05,\ \eta=0.3,\ \Delta=0.1,\ \delta=0.1$ are shown in the left plot of Fig. 12. The effective masses for the so-called principal correlators $\lambda_{0},\lambda_{1}$ are shown, as well as the effective mass corresponding to the off-diagonal correlator $C_{01}$. The principal correlators are the eigenvalues of the matrix $C(t_{0})^{-1/2}\,C(t)\,C(t_{0})^{-1/2}$. Here, we used $t_{0}=5$ with eigenvector pinning to identify the levels. The black curve shows the perfect extraction of $E_{0}$ starting at zero temporal separation from the off-diagonal correlator. The blue curve associated with principal correlator $\lambda_{0}(t)$ eventually plateaus at $E_{0}$, but not until large temporal separations are reached due to the small coupling to the $E_{0}$ state. The red curve is associated with the principal correlator $\lambda_{1}(t)$ which plateaus eventually to $E_{1}$.
This is a contrived and unrealistic (if not impossible to attain) toy model since it has been engineered to produce a perfect energy extraction from the off-diagonal correlator due to the exact zeros in the $Z_{A},Z_{B}$ factors. To achieve this situation in general would require an exact diagonalization of the full volume of lattice operators, not the use of a single operator. Making only a single change, allowing an arbitrarily small, non-zero overlap with these states,
$Z_{A}=(\epsilon_{1},\sqrt{1-\epsilon_{1}^{2}-\epsilon_{2}^{2}},\epsilon_{2}),\quad Z_{B}=(\epsilon_{1},\epsilon_{2},\sqrt{1-\epsilon_{1}^{2}-\epsilon_{2}^{2}}).$ (26)
immediately makes the effect go away, with the off-diagonal correlation function now scaling as,
@@
An even more realistic model is studied in the right-hand plot of Fig. 12. For this plot, two simple changes are made from the original toy model Amarasinghe et al. (2023); Tews et al. (2022): the overlaps with the ground state are not set to be the same for the states created by operators $A$ and $B$, and the exact zeros in the $Z_{A},Z_{B}$ factors are replaced by small but non-zero values:
@@
This model is still unrealistic since there are no couplings to the many higher levels as there would be in actual lattice QCD computations. Lastly, consider a more realistic toy model with 10 levels and 3 operators $A,B,C$ and parameters chosen simply to be
$\mbox{Energies}:(0.2,0.3,0.4,0.45,0.50,0.55,0.60,$
$\qquad\qquad 0.65,0.70,0.75),$
@@
The effective masses corresponding to the three principal correlators $\lambda_{0}(t),\lambda_{1}(t),\lambda_{2}(t)$, as well as the effective mass corresponding to the off-diagonal correlator $C_{01}(t)$, are shown in the top row of Fig. 13. The principal correlators have been determined using $t_{0}=5$ in the left plot, and $t_{0}=25$ in the right plot. The off-diagonal correlator shows no advantage. In the bottom row, a small $\delta=0.005$ is assumed. An advantage is observed in the bottom left plot for $t_{0}=5$, but this goes away as $t_{0}$ is increased, as shown in the bottom right plot.
Figure 13: (Top left) Effective masses corresponding to the three principal correlators $\lambda_{0}(t),\lambda_{1}(t),\lambda_{2}(t)$, shown as blue, red and green curves, respectively, and the effective mass corresponding to the off-diagonal correlator $C_{01}(t)$, shown as a black curve, for the 10-level, 3-operator toy model described in the text, with parameter values given in Eq. (29). Principal correlators have been computed using $t_{0}=5$. (Top right) Same as the plot on the top left, except the principal correlators have been determined using $t_{0}=25$. No advantage is observed for the off-diagonal correlator. (Bottom left and right) Same as top row, except $\delta=0.005$ is used. On the left, an advantage is seen for the off-diagonal correlator, but this goes away as $t_{0}$ is increased (see right).
These toy-models also highlight the importance of carefully choosing the times used in constructing the rotation matrices through the GEVP, particularly $t_{0}$. In the single-pivot method described in Sec. II, the effective masses associated with the diagonal elements of the rotated correlator matrix given in Eq. (3) are guaranteed to monotonically decrease to the lowest-lying energy whose eigenstate has nonzero overlap with the state created by the rotated operator. When the principal correlator method is used (described in Sec. II), as was done to generate Fig. 12 and Fig. 13, this monotonicity is no longer guaranteed since the overlap factors can acquire time dependence. The dramatic variations observed in the effective masses shown Fig. 12 and Fig. 13 for small temporal separations are a symptom of the fact that the operators are poorly coupled to the lowest energy state in the system. In such a situation, the importance of studying the dependence on the extracted spectrum as a function of $t_{0}$ and $t_{d}$ is amplified. The effective energies for two different choices of $t_{0},t_{d}$ using the single-pivot method applied to the toy model of Eq. (29) are shown in Fig. 14. On the left, one sees that $t_{0},t_{d}$ have been chosen much too small, producing a rotated correlator matrix that does not remain diagonal for times larger than $t_{d}$. On the right, the rotated correlator remains diagonal for times larger than $t_{d}$, indicating a good choice of $t_{0},t_{d}$. This great sensitivity to $t_{0},t_{d}$ in this model is again a symptom of the poor operator choices. For the LQCD results presented in this work, we find no evidence of such behavior.
Refer to caption Refer to caption
Figure 14: Effective energies associated with diagonal elements of the rotated correlator matrix of Eq. (3) for the model given in Eq. (29) but with $\delta=0.005$ for two choices of $t_{0},t_{d}$.
Our conclusion is that it is nearly impossible for an off-diagonal correlator to reveal a level hidden from the correlator matrix method, especially for a hexaquark operator that has been shown to contain copious amounts of excited contamination. In fact, negative weights in the spectral representation of an off-diagonal correlator with a slow fall-off of contributions from higher-lying states are more likely to cause plateau misidentifications.
@@
Figure 15: (Left) Comparison of scattering phase shifts in the deuteron channel determined in Ref. Hörz et al. (2021), shown as red, blue, magenta points and magenta and gray bands, with phase shifts determined in Refs. Beane et al. (2013a, b), shown as green points. (Middle/Right) Effective mass, shown as blue points, from Fig. 2 of Ref. Beane et al. (2017) for the deuteron on a $48^{3}$ spatial lattice from an off-diagonal correlator involving a local hexaquark operator at the source and a nucleon-nucleon operator of zero momentum at the sink. The horizontal green line indicates the approximate energy extraction from Fig. 4 of Ref. Beane et al. (2013a), with the violet line indicating approximately where the energy extraction from Ref. Hörz et al. (2021) would be for this $48^{3}$ lattice. (Bottom right) Effective mass, similar to that in the center plot, but for the dineutron.